What mass of zinc is needed to react with 23.1 g of phosphoric acid


1
Expert's answer
2022-02-28T11:04:09-0500

3Zn + 2H3PO4 = Zn3(PO4)2 + 3H2

n = m / M

M (H3PO4) = 98 g/ mol

M (Zn) = 65.4 g/mol


According to the equation, n (Zn) = 3/2 x n (H3PO4)

n (H3PO4) = 23.1 / 98 = 0.24 mol

n (Zn) = 3/2 x n (H3PO4) = 3/2 x 0.24 = 0.36 mol

m (Zn) = n x M = 0.36 x 65.4 = 23.5 g